OpenStudy (anonymous):

A bookcase is to have four shelves, including the top. The height of the bookcase is to be 2 feet more than the width and only 20 feet of lumber is available. Use an equation to determine the width and height of the bookcase. Solve for length and weight. Does any one know how to solve this problem?

OpenStudy (anonymous):

4 shelves, i'll assume that the piece of lumber for the bottom of the bookcase is 1 shelf and the piece at the top is 1 self, so 4w+2h=20 h=w+2 4w+2(w+2)=20 6w+4=20 6w=16 w=16/6 h=16/6+2=14/3